OpenStudy (anonymous):

find all pair integers whose difference is 5 and product is 234

jhonyy9 (jhonyy9):

try solve it this x-y=5 xy=234

OpenStudy (anonymous):

for the first equation can I solve for y or x first. like: x=5/y ?

OpenStudy (anonymous):

Solve for x in the first equation

OpenStudy (anonymous):

\[x=5+y\]

OpenStudy (anonymous):

now i plug that into my second equation

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[y ^{2}+5y-234=0 \]

OpenStudy (anonymous):

now I factor it?

OpenStudy (anonymous):

You got it...

OpenStudy (anonymous):

factoring: (y+18) (y-13) answer: y=-18 and 13

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (anonymous):

Welcome..you knew it all the time